OpenStudy (anonymous):

Calculus II: Is it possible to integrate this question using trig substitution? integral x^3/sqrt(x^2+144) dx

OpenStudy (anonymous):

=Q Q

OpenStudy (anonymous):

sorry i didn't mean to post that

OpenStudy (anonymous):

it's all good

OpenStudy (anonymous):

tangent*

OpenStudy (anonymous):

It's been a long day -- use \[ x = 12 \tan(u) \] \[ dx = 12 \sec^2(u) du\]

OpenStudy (anonymous):

\[\int\frac{x^3}{\sqrt{x^2+144}}dx\] Using @Jemurray3's suggestion, \[\int\frac{(12 \tan u)^3}{\sqrt{(12\tan u)^2+144}}\cdot (12 \sec^2 u \;du)\\ 12^4\int\frac{\tan^3u\sec^2u}{\sqrt{144\tan^2 u+144}}du\\ 12^3\int\frac{\tan^3u\sec^2u}{\sqrt{\tan^2 u+1}}du\] I'm sure you can take it from here.

OpenStudy (anonymous):

I got till \[12^{3}\int\limits_{}^{} \sec ^{3}u \tan u - \tan u \sec (u) du\] Now what to do next?

OpenStudy (anonymous):

Integration by parts?

OpenStudy (anonymous):

Rewrite that as ( something in here ) * sec(u)tan(u) du

OpenStudy (anonymous):

That would be sec^2 (u) - 1

OpenStudy (anonymous):

I think I got it.